# Editable Fairness: Fine-Grained Bias Mitigation in Language Models

### v1

@@ 3.1 Dataset Construction
Create paraphrased social bias dataset. To prevent excessive knowledge retention, we propose to evaluate the generalization ability on further social biases. For social biases in $\Omega_{S}$, we propose generating semantically similar expressions $\Omega_{P}$. To ensure the quality of the generated data, we have conducted meticulous human validation, with details provided in Appendix A.2. We also analyze the diversity and challenge of BiaScope using case examples in Appendix A.3.
3.2 Evaluation Metrics
Retention Score (RS) assesses the percentage of commonsense knowledge in $\Omega_{R}$ retained after debiasing. The evaluation of RS is conducted according to the following criteria:
$\textbf{RS}(\mathcal{G},{\mathcal{G}^{*}},\Omega_{R})=\mathbb{E}_{k_{R}\in\Omega_{R}}\mathbbm{1}\{{\mathcal{G}}[k_{R}]={\mathcal{G}^{*}}[k_{R}]\},$ (1)
where $k_{R}$ represents commonsense knowledge in $\Omega_{R}$. $\mathcal{G}[k_{R}]$ and $\mathcal{G}^{*}[k_{R}]$ represent the prediction of the original and debiased model. $\mathbbm{1}$ is the indicator function.
Paraphrase Stereotype Score (PS) evaluates the generalization ability on paraphrased biases in $\Omega_{P}$. As a complement to RS, it aims to prevent the model from over-retaining knowledge and thereby losing its generalization ability. It computes the percentage of data that a model gives a biased prediction as opposed to an unbiased prediction:
$\textbf{PS}(\mathcal{G}^{*},\Omega_{P})=\mathbb{E}_{k_{p}\in\Omega_{P}}\mathbbm{1}\{\mathcal{P}_{\mathcal{G}^{*}}[k_{p}]>\mathcal{P}_{\mathcal{G}^{*}}[k_{p}^{{}^{\prime}}]\},$ (2)
Figure 3: An illustration of our FAST framework. (a) We first locate the decisive layer towards biased predictions. (b) A fairness stamp is inserted at the decisive layer, which is optimized with the objective of bias mitigation and knowledge retention.
4 Method
Task Formulation.
Given a pre-trained language model $\mathcal{G}$ and a set of social biases $\Omega$ to be calibrated, the task is outputting an edited model $\mathcal{G}^{*}$ where social biases in $\Omega$ can be fairly predicted while other knowledge is retained as in the original $\mathcal{G}$.
Overview.
This section proposes a fine-grained bias mitigation method Fairness-Stamp (FAST). The main idea of FAST is to identify the storage locations of social biases in LLMs and then modify the small number of identified parameters. Thus, it precisely repairs social biases while minimizing the impact on other knowledge. Concretely, FAST operates through a two-step process, as depicted in Figure 3: (1) determining the decisive layer responsible for storing social biases; (2) enveloping the decisive layer with an auxiliary fairness stamp (a 2-layer feed-forward network), which is optimized with the objective of bias mitigation and knowledge retention.
4.1 Step 1: Locate the Decisive Layer for Biases
Definition of Social Bias. In this step, we propose to investigate the storage of social biases in LLMs. Typically, a social bias consists of a certain social group and the biased description, which amplifies social inequalities [69, 4, 1]. For instance, in the statement Black people are more likely to commit a crime,” the phrase are more likely to commit a crime” is the biased description associated with the social group “Black people.” In light of this, we formalize the social bias as a knowledge triplet $k=(s,r,o)$, where $s$ is the subject (i.e., Black people), $o$ is the object (i.e., commit a crime), and $r$ is the relation between them (i.e., are more likely to), inspired by [56, 35].
Contrastive Social Biases Localization. To investigate how social bias $(s_{1},r_{1},o_{1})$ is stored as association between the social group and biased description, we propose to use its counterfactual knowledge $(s_{2},r_{2},o_{2})$ for contrast. This involves altering the social group or biased description (e.g., changing Black people to White people) to better probe these biased associations. Our contrastive bias localization is performed in three runs, with a complete illustration in Figure 6 in the Appendix B.2:
(1) Biased run: We pass the biased prompt $(s_{1},r_{1})$ into the model and collect all hidden states $\{h^{(l)}$ |||| $l\in[1,L]\}$, in a forward run towards biased prediction, where $L$ is number of layers.
(2) Counterfactual run: We pass the counterfactual prompt $(s_{2},r_{2})$ to the model to modify the biased prediction. Hidden states will also change due to the alteration of the input subject.
(3) Restore biased states: To measure the effect of certain layer $\hat{l}$ in the model to the biased prediction, we restore the biased states $h^{(\hat{l})}$ of $s_{1}$ and perform the forward run. Then we calculate the recovery degree of biased prediction, which indicates the effect of layer $\hat{l}$.
Determine the decisive layer. Denote the prediction probability on the object of the biased run as $P[o]$, and the probability of the counterfactual run as $P^{*}[o]$. In this way, the total biased effect (TE) can be defined as: TE $=P[o]-P^{*}[o]$. In the restoration run, the probability will recover from $P^{*}[o]$ to $P[o]$ due to the restoration of certain biased states $h^{(l)}$, which reflects the contribution of these states to the biased prediction. Denote the probability of restoring layer $l$ as $P^{*}(h^{(l)})[o]$. The indirect biased effect (i.e., recovery degree) IE of layer $l$ can be calculated by IE $=P^{*}(h^{(l)})[o]-P^{*}[o]$. The layer with the largest IE will be selected as the decisive layer.
4.2 Step 2: Insert Fairness Stamp at the Decisive Layer
Following Step 1, we propose to select the layer that contributes most significantly to the bias as the decisive layer. Assuming the input hidden states to be $\mathbf{h}$, the decisive layer (i.e., feed-forward network, FFN) in the original LLM can be formulated as follows:
$\text{FFN}(\mathbf{h})=\text{Act}(\mathbf{hK}^{\top})\mathbf{V},$ (3)
where $\mathbf{K}$ and $\mathbf{V}$ denote the parameters (i.e., keys and values matrices) of the first and second linear layers in the FFN, respectively. We propose to envelop the decisive layer with a fairness stamp. The fairness stamp is a 2-layer Feed-Forward Network (FFN) layer, which helps modify the output of the decisive layer with a few external parameters to achieve the goal of fairness. The output of the enveloped FFN layer is given by:
$\text{FFN}^{\prime}(\mathbf{h})=\text{FFN}(\mathbf{h})+\text{Act}(\mathbf{hK^{\prime}}^{\top})\mathbf{V^{\prime}},$ (4)
Bias Mitigation. With a social bias $k_{1}$ and its counterfactual knowledge $k_{2}$, we propose to mitigate the gap between their probabilities of prediction on the objects:
$\mathcal{L}_{e}=\frac{1}{\lvert\Omega\rvert}\sum_{(k_{1},k_{2})\in\Omega}\lvert\mathcal{P}_{\mathcal{G}}[k_{1}]-\mathcal{P}_{\mathcal{G}}[k_{2}]\rvert,$ (5)
where $k_{i}=(s_{i},r_{i},o_{i})$ follows the definition in Section 4.1. $\mathcal{P}_{\mathcal{G}}[k_{i}]=\mathcal{P}_{\mathcal{G}}[o_{i}|p_{i}]=\mathcal{P}_{\mathcal{G}}[o_{i}|s_{i},r_{i}]$ denotes the probability of predicting the object $o_{i}$ given the prompt $p_{i}$.
Knowledge Retention. We propose to retain knowledge in two ways. First, we retain the probability distribution for the input prompts $p_{i}$ to control the deviation from the original model. Second, we retain the probability distribution on the prompt $p^{\prime}$ that combines pre-defined template (e.g., “{subject} is _”) and the input subject (e.g., Black people), which helps retain the perception of different social groups and prevent the model from degradation of knowledge. The two loss functions are as follows:
$\mathcal{L}_{s1}=\frac{1}{\lvert\Omega\rvert}\sum_{p_{i}\in\Omega}\mathcal{D}_{KL}(\mathcal{P}_{\mathcal{G}}[\star|p_{i}],\mathcal{P}_{\mathcal{G}^{*}}[\star|p_{i}]),$ $\mathcal{L}_{s2}=\frac{1}{\lvert\Omega\rvert}\sum_{s_{i}\in\Omega}\mathcal{D}_{KL}(\mathcal{P}_{\mathcal{G}}[\star|p^{\prime}(s_{i})],\mathcal{P}_{\mathcal{G}^{*}}[\star|p^{\prime}(s_{i})]),$ (6)
where $\mathcal{P}_{\mathcal{G}}[\star|p^{\prime}]$ is the predicted probability vector of all objects. $\mathcal{G}$ and $\mathcal{G}^{*}$ represent the origin and debiased model. $\mathcal{D}_{KL}(\cdot,\cdot)$ represents the Kullback-Leibler Divergence.
To prevent the model from overfitting to particular inputs, we utilize prefix texts $x_{j}$ to enhance generalization ability across various contexts. These prefix texts are randomly generated by the model, for instance, “My father told me that”, and are concatenated to the front of the prompts.
The overall objective can be formulated as follows with hyper-parameters $\alpha$ and $\beta$:
5 Experiment
5.1 Experiment details
Models. We employ the representative masked language model BERT (bert-base-uncased[16] and generative language model GPT2 (GPT2-small[57] as our backbones. Extended experiments are conducted on GPT2-XL, GPT-Neo-2.7b [3] and Llama-2-7b [64].
Baselines. We consider the following debiasing techniques as baselines. The techniques can be grouped into two categories. (1) Fine-tuning: Counterfactual Data Augmentation (CDA) [78], Dropout [70], SentenceDebias [43] and Iterative Nullspace Projection (INLP) [58] pre-train, perform dropout on the rebalanced corpus, or remove sensitive attributes from the representations. MABEL [33] mitigates gender bias utilizing a contrastive learning objective on entailment labels. (2) Prompt-tuning: Auto-debias [27] proposes to directly probe the encoded biases through prompts, then mitigate biases via distribution alignment loss. (3) Post-hoc: Self-Debias [61] proposes to leverage internal knowledge to discourage it from generating biased text. FMD [7] proposes a machine unlearning-based strategy to efficiently remove bias. (4) Knowledge Editing: ROME [46] and MEMIT [47] are proposed to effectively and efficiently modify the knowledge in a language model. We adapt their knowledge updating objectives to align with our fairness objectives for comparison, while all other settings remain unchanged.
Datasets and Evaluating Metrics. We evaluate the debiasing performance on StereoSet [52], Crows-Pairs [53], as well as our proposed BiaScope. Stereotype Score (SS), Language Modeling Score (LMS), and Ideal Context Association Test Score (ICAT) correspond to StereoSet to evaluate the extent of bias, the ability to predict reasonable association, and the combined score of the former two, with details in Appendix A.1. SS is also employed by Crows-Pairs to measure bias. As for BiaScope, we utilize RS and PS, as introduced in Section 3.2.
Implementation details. Bias mitigation is conducted over the collected biased knowledge in Section 4.1. We utilize two-layer fully connected neural networks with the ReLU activation function as the fairness stamp, with a hidden dimension of 1024. Additional details are in Appendix C.1.
@@ 5.1 Experiment details
Table 1: Debiasing results on BERT. The best result is indicated in bold. $\diamond$: the closer to 50, the better. “-”: results are not reported. Reported results are means over three training runs. Due to space limitations, results with statistical significance analysis, as well as results in terms of religion are provided in the Appendix C.4.
@@ 5.2 Debiasing Performance
Existing debiasing methods cannot retain individual commonsense knowledge. The debiasing results are delineated in Table 1 and Table 5.2. It is observed that all debiasing baselines fail to yield satisfactory results in knowledge retention (i.e., RS), which proves our claim that group-invariant methods compromise the individual knowledge to distinguish between different social groups.
Our approach surpasses existing debiasing and knowledge-editing baselines in both bias mitigation and knowledge retention. As shown in Table 1 and Table 5.2, our proposed FAST is the first to achieve near-perfect bias mitigation (i.e., SS lower than 52 for BERT) on the two evaluating datasets, while SS of existing approaches, in terms of gender, are still higher than 56. Further, FAST can also largely retain a high RS, and achieve the highest LMS and ICAT. This demonstrates the effectiveness of our fine-grained calibration strategy towards eliminating social biases in LLMs. In addition, we report the performance of knowledge-editing approaches ROME and MEMIT. It can be discerned that neither ROME nor MEMIT significantly improves SS over vanilla BERT. Overall, comparing results demonstrate the effectiveness of our fine-grained calibration strategy towards eliminating social biases in LLMs. Supplemented debiasing results and qualitative studies are in Appendix C.
@@
Our approach scales to larger models. In order to further validate the scalability of FAST, we conduct additional experiments on larger models, i.e., GPT2-XL, GPT-Neo-2.7B, and Llama-2-7B, with results reported in Table 5.2. After debiasing, FAST induces a significant reduction (9.4 in average) in SS, and a great improvement in ICAT. Meanwhile, FAST can also retain the Retention Score for larger language models. These demonstrate the consistent effectiveness and scalability of FAST.
@@ 6 Analysis and Discussion
Language Modeling Capability Analysis.
In this section, we evaluate our debiased models against the General Language Understanding Evaluation (GLUE) benchmark [67] to assess the retention of general linguistic understanding after bias mitigation. As shown in Table 4, FAST achieves better downstream performance than 5 out of 6 baselines on average, indicating that FAST retains language modeling capabilities while mitigating biases.
Method CoLA MNLI MRPC QNLI QQP RTE SST STS-B WNLI Average
BERT 56.78 84.76 89.54 91.51 88.06 64.62 93.35 88.24 56.34 79.24
CDA 2.07 84.84 81.22 84.84 87.85 47.29 92.32 40.83 43.66 62.77
Dropout 2.07 84.78 81.22 91.49 88.02 47.29 92.09 40.87 43.66 63.50
SentDebias 55.72 84.94 88.81 91.54 87.88 63.9 93.12 88.23 56.34 78.94
INLP 56.50 84.78 89.23 91.38 87.94 65.34 92.66 88.73 54.93 77.05
MABEL 57.80 84.50 85.00 91.60 88.10 64.30 92.20 89.20 - -
FAST 55.99 84.75 87.60 91.47 88.12 67.15 92.20 89.05 46.13 78.01
Effectiveness Analysis of Bias Locating.
The locating results of BERT are shown in Figure LABEL:Knowledge_Locating_results_of_BERT.. They reveal that the final layer demonstrates a significantly higher average indirect effect than the other layers, marking it as the decisive layer for bias prediction. To validate the effectiveness of knowledge locating (i.e., Section 4.1), we perform calibration (i.e., step 2) on every layer of BERT, with results shown in Figure LABEL:Effectiveness_of_Knowledge_Locating. It is observable that layer 11 achieves optimal performance in terms of SS, RS, and LMS, corroborating the effectiveness of knowledge locating. Layers 1-5 show minimal alleviation of biases (no decline in SS), suggesting a minimal correlation between these layers with the storage of biased knowledge. Notably, layers 6-10 not only result in a reduction in SS but also a significant decrease in RS, indicating the entanglement of biased knowledge with other knowledge. Locating results of GPT2, GPT2-XL, GPT-Neo-2.7B, and Llama-2-7B are reported in Appendix C.2. We also verify the robustness of knowledge locating in Appendix D.1. Effectiveness validation of fairness stamp (i.e., Section 4.2) is provided in Appendix D.2.
Table 5: Ablation Study on $\alpha$ and $\beta$. Experiments are conducted on BERT in terms of gender. $\diamond$: the closer to 50, the better. The best result is in bold.
@@
Sensitivity Analysis on Hyper-parameters and Fairness-Utility Trade-off Analysis.
We have performed a grid search for hyper-parameters $\alpha$ and $\beta$, with results presented in Table 5. The optimization proves robust within specific ranges (i.e., 20-80 for $\alpha$, 0.05-0.5 for $\beta$). However, a trade-off between the bias mitigation and knowledge retention is observed [37, 44]. When either $\alpha$ or $\beta$ is set to 0, both the knowledge retention score (RS) and language modeling ability (LMS) suffer significant declines. Conversely, when either $\alpha$ or $\beta$ is set too high, the fairness performance (SS) is negatively affected. Based on these findings, we choose $\alpha$ at 40 and $\beta$ at 0.1 as they yield the most favorable overall results.
@@
In this section, we verify the robustness of FAST under limited memory sizes. We alter the dimension of hidden states (dim) in our FAST, thereby changing the number of external parameters. The results are shown in Figure LABEL:Ablation_on_the_Number_of_External_Parameters. The best results are obtained with dim set to 1024. As dim decreases, both SS and RS decline slightly, indicating that a larger number of parameters yields better bias mitigation performance. Further increases in dim do not yield better debiasing results. Therefore, dim is set to 1024. Ablation study on the batch size is provided in Appendix D.6.
@@
In Table 6, we report the number of parameters and operation time of our proposed FAST on the largest and smallest models in our experiments. The time is counted on a single RTX 3090 with one piece of bias. It can be observed that FAST only requires about one percent of parameters and bias mitigation can be finished in less than one second or just a few seconds, indicating that FAST allows for timely debiasing of LLMs.
Table 6: Computational complexity analysis on
BERT and Llama-2. “B” denotes billion.
@@
[…]
GPT-J 26.4 (±0.6) 25.8 (±0.5) 27.0 (±0.5)
FT-W 69.6 (±0.6) 64.8 (±0.6) 24.1 (±0.5)
MEND 19.4 (±0.5) 18.6 (±0.5) 22.4 (±0.5)
ROME 21.0 (±0.7) 19.6 (±0.7) 0.9 (±0.1)
MEMIT 96.7 (±0.3) 89.7 (±0.5) 26.6 (±0.5)
FAST 95.1(±0.4) 90.6 (±0.5) 24.6 (±0.5)
Effectiveness on the Knowledge-editing Task.
We conduct experiments on the knowledge-editing task of Zero-Shot Relation Extraction (zsRE) [40]. We employ GPT-J-6B [68] as backbone, and use baseline methods including the improved Constrained Fine-Tuning (FT+W) [46], MEND [48], ROME [46], and MEMIT [47]. We select layers 3 through 8 for editing, consistent with MEMIT. The training and evaluating datasets are also consistent with MEMIT. As shown in Table 6, our method outperforms most baselines and achieves comparable performance to MEMIT. These findings suggest that our method can be effectively applied to knowledge-editing tasks.
[…]
BERT 60.28 59.17 100.0 84.17
$\checkmark$ - - 47.92 49.27 52.38 66.72
$\checkmark$ $\checkmark$ - 51.76 52.06 92.86 86.93
$\checkmark$ $\checkmark$ $\checkmark$ 51.16 50.80 95.83 86.30
Ablation Study on the Losses.
7 Conclusion and Limitation

### v2

@@ 3.1 Dataset Construction
To prevent excessive knowledge retention, we propose to evaluate the generalization ability on further social biases. For social biases in $\Omega_{S}$, we propose generating semantically similar expressions $\Omega_{P}$. To ensure the quality of the generated data, we have conducted meticulous human validation, with details provided in Appendix A.3. We also analyze the diversity and challenge of BiaScope using case examples in Appendix A.5.
@@
Table 1: Debiasing results on BERT. The best result is indicated in bold. $\diamond$: the closer to 50, the better. “-”: results are not reported. Reported results represent the mean values obtained from three independent training runs. Due to space limitations, results with statistical significance analysis, as well as results in terms of religion are provided in the Appendix C.4.
@@
3.2 Evaluation Metrics
Retention Score (RS)
assesses the percentage of commonsense knowledge in $\Omega_{R}$ retained after debiasing. The evaluation of RS is conducted according to the following criteria:
$\textbf{RS}(\mathcal{G},{\mathcal{G}^{*}},\Omega_{R})=\mathbb{E}_{k_{R}\in\Omega_{R}}\mathbbm{1}\{{\mathcal{G}}[k_{R}]={\mathcal{G}^{*}}[k_{R}]\},$
where $k_{R}$ denotes commonsense knowledge. $\mathcal{G}[k_{R}]$ and $\mathcal{G}^{*}[k_{R}]$ denote the prediction of the original and debiased model. $\mathbbm{1}$ is indicator function.
Paraphrase Stereotype Score (PS)
evaluates the generalization ability on paraphrased biases in $\Omega_{P}$. As a complement to RS, it aims to prevent the model from over-retaining knowledge and thereby losing its generalization ability. It computes the percentage of data that a model gives a biased prediction as opposed to an unbiased prediction:
$\textbf{PS}(\mathcal{G}^{*},\Omega_{P})=\mathbb{E}_{k_{p}\in\Omega_{P}}\mathbbm{1}\{\mathcal{P}_{\mathcal{G}^{*}}[k_{p}]>\mathcal{P}_{\mathcal{G}^{*}}[k_{p}^{{}^{\prime}}]\},$
4 Experiment
4.1 Experiment details
Models.
We mainly employ BERT (bert-base-uncasedDevlin et al. (2018) and GPT2 (GPT2-smallRadford et al. (2019) as our backbones. Extended experiments are conducted on GPT2-XL, GPT-Neo-2.7b Black et al. (2021) and Llama-2-7b Touvron et al. (2023) for scalability.
Baselines.
In this study, we categorize and evaluate debiasing techniques across four main groups: Fine-tuning: Includes Counterfactual Data Augmentation (CDA) Zmigrod et al. (2019), Dropout Webster et al. (2020), SentenceDebias Liang et al. (2020), and Iterative Nullspace Projection (INLP) Ravfogel et al. (2020), focusing on pre-training adjustments and sensitive attribute removal. MABEL He et al. (2022) specifically addresses gender bias using a contrastive learning objective on entailment labels. Prompt-tuning: Auto-debias Guo et al. (2022) uses prompts to probe and mitigate biases through distribution alignment loss. Post-hoc: Self-Debias Schick et al. (2021) leverages internal knowledge to prevent biased text generation, while Fast Model De-biasing (FMD) Chen et al. (2023) employs a machine unlearning strategy to remove bias. Knowledge Editing: ROME Meng et al. (2022a) and MEMIT Meng et al. (2022b) locate and modify model knowledge to align with objectives.
Datasets.
We conduct our experiments on StereoSet Nadeem et al. (2020a) and Crows-Pairs Nangia et al. (2020). StereoSet assesses language models’ propensity to form stereotypes using a fill-in-the-blank challenge. Models select from biased, unbiased, or irrelevant options to complete sentences. CrowS-Pairs features counterfactual sentence pairs that illustrate either biased or unbiased social group associations. We further evaluate on BiaScope (Section 3) for knowledge retention. We also evaluate our debiased models against the General Language Understanding Evaluation (GLUE) benchmark Wang et al. (2018) to assess the general language modeling ability.
Evaluating Metrics.
Stereotype Score (SS) is the most straightforward measure for the bias Nadeem et al. (2020b); Nangia et al. (2020). It computes the percentage of knowledge for which a model assigns the biased object as opposed to the unbiased object. Language Modeling Score (LMS), employed in StereoSet Nadeem et al. (2020b), represents the percentage that a model that prefers a relevant association (either the biased object or the unbiased object) as opposed to an irrelevant object. Ideal Context Association Test Score (ICATNadeem et al. (2020a) combines both LMS and SS by $\text{ICAT}=\text{LMS}*\text{min}(\text{SS},100-\text{SS})/50$. It represents the language modeling ability of a model while behaving in an unbiased manner. As for BiaScope, we utilize RS and PS, as in Section 3.2.
Implementation details.
We utilize two-layer fully connected neural networks with the ReLU activation function as the fairness stamp, with a hidden dimension of 1024. We use Adam optimizer with a learning rate of 0.1. We train each batch for 20 iterations. $\alpha$ is set to be 40 and $\beta$ is 0.1. Additional details are in Appendix C.1.
Method CoLA MNLI MRPC QNLI QQP RTE SST STS-B WNLI Average
BERT 56.78 84.76 89.54 91.51 88.06 64.62 93.35 88.24 56.34 79.24
CDA 2.07 84.84 81.22 84.84 87.85 47.29 92.32 40.83 43.66 62.77
Dropout 2.07 84.78 81.22 91.49 88.02 47.29 92.09 40.87 43.66 63.50
INLP 56.50 84.78 89.23 91.38 87.94 65.34 92.66 88.73 54.93 77.05
MABEL 57.80 84.50 85.00 91.60 88.10 64.30 92.20 89.20 - -
FAST 55.99 84.75 87.60 91.47 88.12 67.15 92.20 89.05 46.13 78.01
@@ 4.2 Debiasing Performance
Existing debiasing methods cannot retain individual commonsense knowledge.
The debiasing results are delineated in Table 1 and Table 3. It is observed that all debiasing baselines fail to yield satisfactory results in knowledge retention (i.e., RS), which proves our claim that group-invariant methods compromise the individual knowledge to distinguish between different social groups.
Our approach surpasses baselines in both bias mitigation and knowledge retention.
As shown in Table 1 and Table 3, our proposed FAST is the first to achieve near-perfect bias mitigation (i.e., SS lower than 52 for BERT) on the two evaluating datasets, while SS of existing approaches, in terms of gender, are still higher than 56. Further, FAST can also largely retain a high RS, and achieve the highest LMS and ICAT. This demonstrates the effectiveness of our fine-grained calibration strategy towards eliminating social biases in PLMs. In addition, we report the performance of knowledge-editing approaches ROME and MEMIT. It can be discerned that neither ROME nor MEMIT significantly improves SS over vanilla BERT. Overall, comparing results demonstrate the effectiveness of our fine-grained calibration strategy towards eliminating social biases in PLMs. Supplemented debiasing results are in Appendix C.
@@
Our approach scales to larger models.
In order to further validate the scalability of FAST, we conduct additional experiments on larger models, i.e., GPT2-XL, GPT-Neo-2.7B, and Llama-2-7B, with results reported in Table 4. After debiasing, FAST induces a significant reduction (9.4 in average) in SS, and a great improvement in ICAT. Meanwhile, FAST can also retain the Retention Score for larger language models. These demonstrate the consistent effectiveness and scalability of FAST.
@@
Our approach retains language modeling capability while mitigating bias.
As shown in Table 2, FAST achieves better downstream performance than 5 out of 6 baselines on average, indicating that FAST retains language modeling capabilities while mitigating biases. In summary, these results substantiate that FAST addresses the proposed issue in existing methods where the pursuit of equity compromises the preservation of other existing knowledge. Moreover, empirical evidence confirms the effectiveness of our localize-and-mitigate framework in identifying and mitigating specific biased knowledge, thereby validating Assumption 1.
@@ 5 Analysis and Discussion
Language Models as Social Bias Knowledge Bases.
In our experiments, we select the last layer of BERT as the decisive layer as it demonstrates a significantly higher average indirect effect than the other layers, as shown in Figure LABEL:Knowledge_Locating_results_of_BERT.. To confirm that bias social knowledge are indeed stored in the localized decisive layer, we perform FAST on every layer of BERT, with results shown in Figure LABEL:Effectiveness_of_Knowledge_Locating. It is observable that layer 11 achieves optimal performance in terms of SS, RS, and ICAT, corroborating the effectiveness of knowledge locating. Layers 1-5 show minimal alleviation of biases (no decline in SS), suggesting a minimal correlation between these layers with the storage of biased knowledge. Notably, layers 6-10 not only result in a reduction in SS but also a significant decrease in RS, indicating the entanglement of biased knowledge with other knowledge. This suggests that our framework can identify where social bias knowledge is stored in language models. Additional results and analysis can be referred to Appendix C.2 and D.3.
Fairness-Utility Trade-off via Hyperparameters.
We have performed a grid search for hyperparameters $\alpha$ and $\beta$, with results presented in Table 5. The optimization proves robust within specific ranges (i.e., 20-80 for $\alpha$, 0.05-0.5 for $\beta$). However, a trade-off between the bias mitigation and knowledge retention is observed Kim et al. (2020); Liu and Vicente (2022). When either $\alpha$ or $\beta$ is set to 0, both the knowledge retention score (RS) and language modeling ability (LMS) suffer significant declines. Conversely, when either $\alpha$ or $\beta$ is set too high, the fairness performance (SS) is negatively affected. Based on these findings, we choose $\alpha$ at 40 and $\beta$ at 0.1 as they yield the best overall results.
Table 5: Sensitivity Analysis on $\alpha$ and $\beta$. Experiments are conducted on BERT in terms of gender. $\diamond$: the closer to 50, the better. The best result is in bold.
@@
In this section, we evaluate the robustness of the FAST framework by varying the dimension of hidden states (dim), impacting the number of external parameters. Results, shown in Figure LABEL:Ablation_on_the_Number_of_External_Parameters, indicate optimal performance at dim = 1024. Reduction in dim leads to a slight decrease in SS and RS metrics, supporting the advantage of higher parameter counts for enhanced bias mitigation. No additional benefits are observed with dim increments beyond 1024. Thus, we set dim to 1024 for balance. Details on batch size effects are discussed in Appendix D.4.
@@
In Table 6, we present the parameter count and average processing time for a single social bias case using our proposed FAST framework on both the largest and smallest models tested in our experiments. These measurements were taken on a single RTX 3090. It is evident that FAST requires only about one percent of the parameters and can complete bias mitigation in under one second or just a few seconds. This demonstrates that FAST enables lightweight and efficient debiasing in PLMs.
Table 6: Computational complexity analysis on BERT and Llama-2. “B” denotes billion.
@@
6 Conclusion